# STACKFEED: Structured Textual Actor-Critic Knowledge base editing with FEEDback

### v1

@@ 6.2.1 Incomplete Knowledge Base
We adapt two code generation datasets from ARKS (Su et al., 2024a), namely ARKS-Pony and ARKS-Ring. The dataset consists of LeetCode problems and their solutions in low-resource languages Pony and Ring respectively. Each datapoint is supplemented with a corresponding language documentation, with execution accuracy as the success metric and execution failures as feedback to the system. Given that these language don’t appear prominently in LLM pre-training data, the performance of code generation RAG agents on these datasets depends significantly on the quality of the Knowledge Base. However, given that these languages have smaller communities, their documentation isn’t as well maintained and often lack critical information. . For the purpose of evaluation on these datasets, we split them into train, eval, test splits as specified in Table  1. To ensure that we have a good representation of failure cases during training, we first execute the RAG pipeline on the entire dataset and divide the failures at random in a 1:1:2 ratio for train, eval and test respectively. All the datapoints with successful execution match are put in the test split. We use the compiler feedback from the executions as the expert feedback to the STACKFEED system.
6.2.2 Incorrect Knowledge Base
@@
For evaluating under this setting, we leverage the ARKS-ScipyM and ARKS-TensorflowM datasets from ARKS and the CLARK-news dataset from Erase (Li et al., 2024). The ARKS datasets consist of data science problems sourced from the DS-1000 dataset  (Lai et al., 2022), which are to be solved by artificially perturbed versions of scipy and tensorflow libraries respectively, while referring to the original unperturbed documentation. Similar to Pony and Ring, we use the execution accuracy on a test bench as a success metric and use compiler outcome as expert feedback. We also follow a similar approach for data splitting.
Dataset Ring Pony SciPy Tensorflow CLARK-news
Acc $\sigma$ Acc $\sigma$ Acc $\sigma$ Acc $\sigma$ Acc $\sigma$
Base KB 30.77 2.09 29.99 1.57 52.04 0.00 28.88 2.18 26.27 1.20
PromptAgent-E 33.33 2.81 32.22 1.57 53.40 3.12 47.77 3.57 28.80 2.39
STACKFEED 36.75 1.21 37.04 1.28 59.38 1.22 53.84 3.11 37.28 1.69
Table 2: Generalization performance comparison between STACKFEED and baseline models across multiple datasets, reported as accuracy percentages (higher is better).
@@
While fact retrieval is one of the most popular use cases of RAG systems, evolving nature of information requires us to keep the knowledge bases up to date. To simulate these dynamic factual knowledge updates we use the CLARKS-news dataset from Erase (Li et al., 2024) which contains questions and their respective answers extracted from Wikidata at different timestamps. Each timestamp is characterized by a set of articles that were added in the data at that time. For our evaluation, we pool all the questions whose answers changed for the first time at a given timestamp and split them across train, eval and test splits in a 1:1:2 ratio (Table 1).
6.3 Evaluation Metrics
In section 5 we discussed the desirable properties of a Knowledge Base edit. We leverage these properties to design 3 metrics for the KB Editing problem as follows:
Completeness: We use the train set accuracy to estimate the degree of expert feedback incorporated in the learnt Knowledge Base.
Generalization: To estimate the degree of generalization of our Knowledge Base edits, we use the held out test set accuracy.
Coherence: To quantify the degree of coherence of the KB, we first calculate a document-wise coherence score using G-Eval (Liu et al., 2023) with GPT4-1106-preview  as the judge model. The G-eval prompt assigns a 1-5 score to the diff of changes with respect to the original document, checking for thematic similarity of the diff. We pool all the edited documents for a KB edit and average there respective coherence score to define the KB coherence metric.
6.4 System Configurations
MCTS parameters: We use the Upper Confidence bounds applied to Trees (UCT) algorithm for selecting expansion nodes, enabling effective exploration and exploitation of the KB state space. For our experiments, we set a maximum search depth of $3$, an expansion width of $3$, and a maximum of $5$ iterations. The UCT exploration constant is set to $2.5$. These parameters were chosen to balance the computational cost and the need for adequate exploration. A depth of $3$ ensures that the search can explore sufficient variations in the KB states without unnecessary expansion, while an expansion width of $3$ allows a moderate number of candidate states to be evaluated at each step. Similarly, $5$ iterations provide enough opportunity to refine the state search, and the UCT constant of $2.5$ encourages sufficient exploration in early stages while converging towards high-reward states in later stages. For unstructured data, the documents are chunked after every 50 lines and then edit the chunks.
RAG System: For the purpose of our evaluations, we setup a generic RAG system which uses an embedding similarity for semantic retrieval. Additionally, in lines with prior works like  (Zhang et al., 2023) for coding related tasks, we use an iterative retrieval setup wherein we first generate a code using naive retrieval and then query the database again with both the question and generated code to improve the quality of retrieval before generating the final result.
LLM configs: We use OpenAI-Text-Embedding-3-Large  as the embedding model with dimensions size of 3072 and use cosine similarity as a metric of embedding match for ranking. To account for the 8191 max input limit, we create document chunks of at most 7500 tokens. For the reasoning model, we use GPT4-1106-preview, with a temperature of 00. Since LLMs are known to perform poorly with longer context input (Liu et al., 2024), we restrict the max token budget for retrievals at 18000 tokens and remove any lower ranked retrieval to fit this token budget.
7 Results
Completeness( in %) Coherence (1-5, higher is better)
Dataset Ring Pony SciPy Tensorflow CLARK-news Ring Pony SciPy Tensorflow CLARK-news
PromptAgent-E 4.27 3.22 33.33 33.33 11.86 4.33 1.86 2.0 4.0 1
STACKFEED 8.98 9.68 31.38 44.44 13.79 4.67 4.6 4.30 4.0 1
7.1 Completeness and Generalization
We observe consistent improvements over the PromptAgent-E baseline in completeness and generalizability scores, with STACKFEED achieving approximately 2x performance gains on Ring and Pony datasets. However, feedback incorporation remains limited, likely due to suboptimal retrieval or limited document-query associations hindering learnability of edit. STACKFEED also demonstrates higher generalizability and lower variance, attributed to its structured and focused document edits that enhance coherence.
7.2 STACKFEED makes high quality coherent edits
As seen in Table 3, STACKFEED produces edits with a coherence score of 4 or higher for most datasets. For KBs which need long term maintenance (like language and code documentation as seen in the ARKS datasets), STACKFEED makes more coherent edits compared to the baseline. This is especially true for long documents as seen in the ARKS Pony dataset. For news-article like dataset like CLARK-news with factual edits. Incoherency is naturally induced when the facts of the article are changed. For instance, an article on the coronation of a king will lose coherency when the article is updated to add information about the coronation of a new king.
8 Conclusion
We introduced STACKFEED, a novel framework for refining Knowledge Bases (KBs) in Retrieval-Augmented Generation (RAG) systems using a multi-actor, centralized critic architecture. STACKFEED enables efficient KB updates without retraining or altering model parameters by leveraging feedback-driven structured edits and textual gradients.
Our approach achieved superior performance in preserving knowledge base (KB) coherence, consistency, and completeness, resulting in enhanced RAG system responses. Nonetheless, there remains considerable potential for further advancements. Future work will focus on refining these three metrics to elevate system performance even further.
Appendix A Appendix
A.1 Prompts used in STACKFEED
@@
""" There exists a Language Model based software named CodeRAG that automatically does the following task for a developer: task - task_desc
CodeRAG uses a knowledge base to perform this task: kb_desc
Here is one knowledge base file that was involved in these errors: """ for i, file in enumerate(kb_files): prompt += f""" File i+1: id: file[’id’] content:
n<file>
nfile[’content’]
n</file>
n""" if "special_notes" in file and file["special_notes"] != "": prompt += f"""
nspecial_notes: file[’special_notes’]"""
""" The following are the reflections on the errors made by CodeRAG: reflections_str
The reflections show the relationship of the file with the errors made by CodeRAG. If the file is named "None," it means the information about the error on which the reflection is based does not fully fit any knowledge base file.
@@
You have a scratchpad to reason and plan your generalization. Your scratchpad is for your use only and will not be shared with anyone else. The scratchpad is represented by the <scratchpad></scratchpad> tags.
Your generalization should follow this format: <scratchpad> The contents of the scratchpad </scratchpad> <generalization> Your generalization for this file </generalization>
@@
General guidelines: 1. Carefully analyze the reflections to understand the errors CodeRAG is making. 2. "None" is a special file, representing that to fix the error, the information should be in a new file. """
Figure 3: Generalization Stage Prompt
[…]
""" Your task is to reflect upon the errors made by CodeRAG based on the user feedback and provide a reflection on the role of the knowledge base files in the making of those errors.
@@
    You should first analyze the question, the test bench, the feedback, and the output to understand the error made by CodeRAG.
  2. 2.
@@
    Then you should carefully analyze the knowledge base files to see if the theme and the contents of any knowledge base file are relevant to the error. Particularly, you should look out for files that have a factual error related to the error or are missing some information which should have been in the file according to the theme of the file.
    1. (a)
@@
      Read the content of the file and understand the theme of the file. The theme of this file is of course based on the file ID and the content of the file but you should also consider its positioning in the knowledge base. That means you should consider the other files that were extracted for the code file and see how this file fits in with them. For example, if the file is a very basic general guide to the task with other files providing more detailed information, then it would make sense for this file to not have detailed information about specific cases.
    2. (b)
@@
      See if the file has any information related to the error. Check for relevant keywords and how the file might have biased the language model to make the error.
    3. (c)
@@
      If the file has information related to the error, see if the information is correct and complete. If the information is incorrect or incomplete, the file is responsible for the error.
    4. (d)
@@
      If it doesnt́ have information related to the error, check if it makes sense for the file to have information related to the error. If it doesn’t make sense, the file is not responsible for the error. When deciding this, check whether the information would be better suited in any of the other knowledge base files. If the missing info fits better in another file, then deem this file to not be responsible for the error as the missing content can be better placed in the other file.
    5. (e)
@@
      If the file is responsible for the error, explain the error in your reflection and set the needs_editing flag to True. And if the file is not responsible for the error, set the needs_editing flag to False.
  3. 3.
@@
    If none of the files have any error or if you think the content for the error should be in a new file, put a file with the name ‘‘None’’ in your reflection and for its reflection, describe the error and mention why it is not due to the knowledge base files. For the ‘‘None’’ file, the needs_editing flag should always be set to True. The ‘‘None’’ file should be placed as File n+1 where n is the number of files extracted for the code file.
  4. 4.
@@
    Choose the least number of files for editing, we want to change as few files as we can for any error. For example, if we have 5 knowledge base files, unless very extreme cases, we wouldn’t want to set the needs_editing flag as True on more than 2 files. Figure out what the most relevant files for the error are and focus on them.
  5. 5.
@@
    When you choose to edit multiple files, you should make sure that their involvements in the error are distinct and not overlapping. If they are overlapping, think about whether changing one file would be enough to fix the error.
Figure 4: Selection Stage Prompt
[…]
@@
""" There exists a Language Model based software named CodeRAG that automatically does the following task for a developer:
@@
Figure 5: Reflection Stage Prompt
[…]
There exists a Language Model based software named CodeRAG that automatically does the following task
for a developer:
{task} - {task_desc}
CodeRAG was used to perform the task on a specific repository.
The following files in the repository were edited by CodeRAG:
<files>
{files}
</files>
A developer has provided the following feedback on CodeRAG’s output:
<feedback>
{feedback}
</feedback>
Your task is to parse this feedback.
You must separate the larger feedback into smaller feedbacks, where each feedback corresponds
to a specific file in the repository.
**DO NOT** change the content of the feedback. Your job is only to split the feedback into file-level
feedbacks, without altering the feedback’s content in any way.
You should respond in the following output format:
‘‘‘json
    "task": The task name, like {task},
    "num_files": The number of files mentioned in the feedback,
    "feedback_files": [
            "target_file": The file name that the feedback is about, it should be the
            whole path of the file, leave it empty if you are not sure which file the feedback is about. If
            the file name is not empty,
            it should be one of the files in the repository.
            If the file name does not exist in the repository, put an empty string here,
            "num_feedbacks": The number of feedbacks for this file,
            "feedbacks": [
                    "feedback_tone": The tone of the feedback, can take the values ‘positive‘ or
                    ‘negative‘,
                    "target_spans": The spans of the code that the feedback is about.
                    It is a list of json objects with two keys, ’start’ and ’end’.
                    Put an empty list if you are not sure which code span the feedback is about,
                    "feedback": Description of the feedback,
Make sure to include the backticks (‘‘‘) surrounding the output. They are needed for further parsing of your output. """
Figure 6: Feedback Parsing Prompt

### v2

@@ A.5.1 Incomplete Knowledge Base
We adapt two code generation datasets from Evor (Su et al., 2024), namely Evor-Pony. The dataset consists of LeetCode problems and their solutions in low-resource languages Pony and Ring respectively. Each datapoint is supplemented with a corresponding language documentation, with execution accuracy as the success metric and execution failures as feedback to the system. Given that these languages don’t appear prominently in LLM pre-training data, the performance of code generation RAG agents on these datasets depends significantly on the quality of the Knowledge Base. However, given that these languages have smaller communities, their documentation isn’t as well maintained and often lacks critical information. . For the purpose of evaluation on these datasets, we split them into train, eval, and test splits as specified in Table 5. To ensure that we have a good representation of failure cases during training, we first execute the RAG pipeline on the entire dataset and divide the failures at random in a 1:1:2 ratio for train, eval, and test respectively. All the datapoints with successful execution matches are put in the test split. We use the compiler feedback from the executions as the expert feedback to the STACKFEED system.
A.5.2 Incorrect Knowledge Base
@@
For evaluating under this setting, we leverage the Evor-ScipyM and Evor-TensorflowM datasets from Evor and the CLARK-news dataset from Erase (Li et al., 2025a). The Evor datasets consist of data science problems sourced from the DS-1000 dataset (Lai et al., 2022), which are to be solved by artificially perturbed versions of scipy and tensorflow libraries respectively, while referring to the original unperturbed documentation. Similar to Pony and Ring, we use the execution accuracy on a test bench as a success metric and use compiler outcomes as expert feedback. We also follow a similar approach for data splitting.
@@
While fact retrieval is one of the most popular use cases of RAG systems, evolving nature of information requires us to keep the knowledge bases up to date. To simulate these dynamic factual knowledge updates we use the CLARK-news dataset from Erase (Li et al., 2025a) which contains questions and their respective answers extracted from Wikidata at different timestamps. Each timestamp is characterized by a set of articles that were added in the data at that time. For our evaluation, we pool all the questions whose answers changed for the first time at a given timestamp and split them across train, eval and test splits in a 1:1:2 ratio (Table 5).
A.6 Example of Automatic Edits in Evor-Scipy
This is a case in which both PromptAgent-E and STACKFEED opt to append sections to the end of a document about sparse matrices in the modified version of SciPy provided in the Evor Dataset.
Demonstrably, the baseline edits add unnecessary information from the newScience.algorithm and newScience.Distribution modules to the document. The document is about sparse matrices so the addition of information about lineSearch and norm_ppf is not appropriate for this document and it is causing the document to become incoherent
On the other hand, the edits made by STACKFEED are relevant to sparse matrices and keep the document coherent.
Listing 2: PromptAgent-E Edits (Only showing the added sections)
+# newScience.algorithm Module
+## Functions
* lineSearch(func, grad, initial_point, direction, **kwargs)
Perform a line search to find the step size (alpha)
that satisfies the strong Wolfe conditions.
Parameters:
- func : callable
The objective function to be minimized.
- grad : callable
The gradient of the objective function.
- initial_point : ndarray
The starting point for the line search.
Must be provided as a NumPy array.
# newScience.distribution Module
## Functions
* norm_ppf(p_values)
Convert p-values to z-scores using the standard normal
distribution. This function is also known as the inverse
of the cumulative distribution function (CDF) for the
standard normal distribution.
Parameters:
- p_values : ndarray
An array of p-values to be converted.
Returns:
- z_scores : ndarray
The corresponding z-scores for the given p-values.
Example:
‘‘‘python
import newScience.distribution as nsd
import numpy as np
p_values = np.array([0.05, 0.025, 0.975])
z_scores = nsd.norm_ppf(p_values)
print(z_scores)
Listing 3: STACKFEED Edits (Only showing added sections)
### Custom Sparse Matrix Classes
When working with custom implementations of sparse matrices,
it is crucial to understand that the attributes and methods
described in this documentation are specific to the
scipy.sparse.csr_matrix class.
Custom sparse matrix classes may not have the same interface or
support the same operations.
#### Disclaimer on Custom Implementations
Developers should not assume that custom sparse matrix classes
will have attributes like nnz or methods such as nonzero’.
Instead, they should refer to the documentation or
implementation of the custom class to understand the available
attributes and methods.
##### Example of Incorrect Assumption
‘‘‘python
# Incorrect assumption that a custom SparseMatrix class has
nnz attribute
sparse_matrix = SparseMatrix(...) # Custom sparse matrix class
if sparse_matrix.nnz == 0:
print("The sparse matrix is empty.")
# This may raise AttributeError if nnz is not an attribute of
SparseMatrix
Example of Correct Usage
python
Copy
Edit
# Correct approach using methods supported by the custom class
sparse_matrix = SparseMatrix(...) # Custom sparse matrix class
if sparse_matrix.count_nonzero() == 0:
print("The sparse matrix is empty.")
# Alternatively, convert to a dense array and use numpys
count_nonzero
dense_array = sparse_matrix.toarray()
if np.count_nonzero(dense_array) == 0:
print("The sparse matrix is empty.")
A.7 Prompts used in STACKFEED
There exists a Language Model based software named CodeRAG that automatically does the following task for a developer:
{task} - {task_desc}
CodeRAG uses a knowledge base to perform this task:
{kb_desc}
Here is one knowledge base file that was involved in these errors:
for i, file in enumerate(kb_files):
prompt += f"""
File {i+1}:
id: {file[’id’]}
@@ A.7 Prompts used in STACKFEED
if "special_notes" in file and file["special_notes"] != "":
[…]
The following are the reflections on the errors made by CodeRAG:
{reflections_str}
The reflections show the relationship of the file with the errors made by CodeRAG.
If the file is named "None," it means the information about the error on which the reflection is based does not fully fit any knowledge base file.
@@
You have a scratchpad to reason and plan your generalization. Your scratchpad is for your use only and will not be shared with anyone else.
The scratchpad is represented by the <scratchpad></scratchpad> tags.
Your generalization should follow this format:
<scratchpad>
The contents of the scratchpad
</scratchpad>
<generalization>
Your generalization for this file
</generalization>
@@
General guidelines:
1. Carefully analyze the reflections to understand the errors CodeRAG is making.
2. "None" is a special file, representing that to fix the error, the information should be in a new file.
Listing 4: Generalization Stage Prompt
Your task is to reflect upon the errors made by CodeRAG based on the user feedback and provide a reflection on the role of the knowledge base files in the making of those errors.
@@
1. You should first analyze the question, the test bench, the feedback, and the output to understand the error made by CodeRAG.
@@
2. Then you should carefully analyze the knowledge base files to see if the theme and the contents of any knowledge base file are relevant to the error. Particularly, you should look out for files that have a factual error related to the error or are missing some information which should have been in the file according to the theme of the file.
@@
a. Read the content of the file and understand the theme of the file. The theme of this file is of course based on the file ID and the content of the file but you should also consider its positioning in the knowledge base. That means you should consider the other files that were extracted for the code file and see how this file fits in with them. For example, if the file is a very basic general guide to the task with other files providing more detailed information, then it would make sense for this file to not have detailed information about specific cases.
@@
b. See if the file has any information related to the error. Check for relevant keywords and how the file might have biased the language model to make the error.
@@
c. If the file has information related to the error, see if the information is correct and complete. If the information is incorrect or incomplete, the file is responsible for the error.
@@
d. If it doesnt have information related to the error, check if it makes sense for the file to have information related to the error. If it doesnt make sense, the file is not responsible for the error. When deciding this, check whether the information would be better suited in any of the other knowledge base files. If the missing info fits better in another file, then deem this file to not be responsible for the error as the missing content can be better placed in the other file.
@@
e. If the file is responsible for the error, explain the error in your reflection and set the needs_editing flag to True. And if the file is not responsible for the error, set the needs_editing flag to False.
@@
3. If none of the files have any error or if you think the content for the error should be in a new file, put a file with the name "None" in your reflection and for its reflection, describe the error and mention why it is not due to the knowledge base files. For the "None" file, the needs_editing flag should always be set to True. The "None" file should be placed as File n+1 where n is the number of files extracted for the code file.
@@
4. Choose the least number of files for editing, we want to change as few files as we can for any error. For example, if we have 5 knowledge base files, unless very extreme cases, we wouldnt want to set the needs_editing flag as True on more than 2 files. Figure out what the most relevant files for the error are and focus on them.
@@
5. When you choose to edit multiple files, you should make sure that their involvements in the error are distinct and not overlapping. If they are overlapping, think about whether changing one file would be enough to fix the error.
Listing 5: Selection Stage Prompt
@@
There exists a Language Model based software named CodeRAG that automatically does the following task for a developer:
@@
Listing 6: Reflection Stage Prompt